OpenStudy (anonymous):

Six multiplied by the reciprocal of a number is equal to 2/3. Find the number.

OpenStudy (dumbcow):

6*(1/n) = 2/3 6/n = 2/3 cross-multiply 2n = 18 n = 9

OpenStudy (anonymous):

Thank you once again.